An airplane flew for 6 hours with a tail of 60 km/hr. The return flight against the same wind took 8 hours. Find the speed of the airplane in still air.
Answer by stanbon(48558)

An airplane flew for 6 hours with a tail of 60 km/hr. The return flight against the same wind took 8 hours. Find the speed of the airplane in still air.
-----------
With the wind DATA:
Time = 6 hrs ; rate = p+60 mph ; distance = tr = 6(p+60) = 6p+360 miles
---------
Against the wind DATA:
Time = 8hrs ; rate = p-60 mph ; distance = tr = 8(p-60) = 8p-480 miles
---------
EQUATION:
distance with = distance against
6p+360 = 8p-480
2p = 840
p = 480 mph (speed of the plane in still air)
===========
Cheers,
Stan H.